OpenStudy (anonymous):

Graph the following quadratic functions, locating the vertices and x- intercepts (if any) accurately.

OpenStudy (anonymous):

\[f (x) = (x-2)^2 +3\]

OpenStudy (anonymous):

if the parabola is written as f(x)=a(x-h)^2+k the vertex is (h,k) in your case, vertex is (2, 3)

OpenStudy (anonymous):

to find the x-intercepts, expand (x-2)^2+3 to x^2-4x+7 and set it to 0 x^2-4x+7=0 solve for x since the discriminant, b^2-4ac<0 in this case, there are no x-intercepts because the solution to this quadratic equation has 2 complex roots

OpenStudy (anonymous):

Wait i got confused? How do i graph it?

OpenStudy (anonymous):

do you know how to graph y=x^2?

OpenStudy (anonymous):

I think so but i am not sure..

OpenStudy (anonymous):

if you can graph y=x^2, then you can graph y=(x-2)^2+3 y=(x-2)^2+3 is the graph of y=x^2 shifted right 2 units and up 3 units

OpenStudy (anonymous):

I wish you could draw graphs on here.

OpenStudy (anonymous):

Can someone please show me how to graph this?

OpenStudy (apoorvk):

yes you can. first find the vertex of the curve. its (-b/2a , -D/4a) so that makes it (2,-3). next find the roots. they are none ofcourse. so, curve a symmetric parabola with vertex (2,3). and find out a few cordinates (like y-coordinates) for accuracyof the plot (ofcourse it will be free hand)= i hope you get an idea (i am sorry the drawing couldnt be better)